# RL-MILP Solver: A Reinforcement Learning Approach for Solving Mixed-Integer Linear Programs with Graph Neural Networks

### v1

@@ Abstract
Mixed-Integer Linear Programming (MILP) is an optimization technique widely used in various fields. Primal heuristics, which reduce the search space of MILP, have enabled traditional solvers (e.g., Gurobi) to efficiently find high-quality solutions. However, traditional primal heuristics rely on expert knowledge, motivating the advent of machine learning (ML)-based primal heuristics that learn repetitive patterns in MILP. Nonetheless, existing ML-based primal heuristics do not guarantee solution feasibility (i.e., satisfying all constraints) and primarily focus on prediction for binary decision variables. When addressing MILP involving non-binary integer variables using ML-based approaches, feasibility issues can become even more pronounced. Since finding an optimal solution requires satisfying all constraints, addressing feasibility is critical. To overcome these limitations, we propose a novel reinforcement learning (RL)-based solver that interacts with MILP to find feasible solutions, rather than delegating sub-problems to traditional solvers. We design reward functions tailored for MILP, which enables the RL agent to learn relationships between decision variables and constraints. Additionally, to effectively model complex relationships among decision variables, we leverage a Transformer encoder-based graph neural network (GNN). Our experimental results demonstrate that the proposed method can solve MILP problems and find near-optimal solutions without delegating the remainder to traditional solvers. The proposed method provides a meaningful step forward as an initial study in solving MILP problems end-to-end based solely on ML.
@@ 1 Introduction
The Traveling Salesman Problem and the Knapsack Problem are representative examples of combinatorial optimization (CO) problems that have been extensively studied in operations research and computer science. CO addresses mathematical optimization that aim to minimize or maximize the value of a specific objective function. If the objective function and constraints of CO are linear, it is called linear programming (LP). If some decision variables in LP must take integer values, it becomes Mixed-Integer Linear Programming (MILP) (Bengio, Lodi, and Prouvost 2021). MILP is widely used to model various problems (Zhang et al. 2023) such as logistics (Kweon et al. 2024), path planning (Zuo et al. 2020), and energy systems (Ren and Gao 2010).
Primal heuristics aim to quickly find initial solutions to MILP by reducing the search space (Achterberg, Berthold, and Hendel 2012). Traditional primal heuristics are crafted based on experts’ knowledge. However, this approach requires significant design effort and poses challenges for practitioners lacking domain expertise (Bengio, Lodi, and Prouvost 2021). In contrast, recent advancements have introduced ML-based primal heuristics (Gasse et al. 2019; Nair et al. 2020; Yoon 2022; Han et al. 2023; Cantürk et al. 2024). These approaches have gained attention because MILP instances generated from a specific distribution often share similar patterns, and ML is effective at learning such patterns (Gasse et al. 2022). Figure 1(a) illustrates how ML-based primal heuristics solve MILP. A trained ML model generates a partial solution for some integer variables in a given MILP instance. The generated partial solution is then passed to a traditional optimizer (e.g., Gurobi), which optimizes the remaining sub-problem to derive the final solution.
@@
Figure 1: Process of solving MILP using ML.
@@
Existing ML-based primal heuristics have demonstrated their ability to quickly find good solutions by reducing the problem space. However, inaccurate ML predictions hinder solution feasibility (i.e., satisfying all constraints), emphasizing the need to consider feasibility (Han et al. 2023). The infeasibility prevents reaching the optimal solution, posing a significant obstacle to solving MILP. Moreover, most ML-based primal heuristics (Gasse et al. 2019; Yoon 2022; Han et al. 2023; Cantürk et al. 2024) focus solely on predicting values for binary decision variables. However, many real-world problems involve non-binary integer (integer, for short) variables (e.g., logistics (Kweon et al. 2024), maritime transportation (Papageorgiou et al. 2014), and energy systems (Ren and Gao 2010)). This highlights the need for techniques capable of effectively handling integer variables.
@@
Infeasibility caused by incorrect ML predictions can be more pronounced for integer variables due to their broader prediction range than binary variables. For instance, imagine a naive ML approach that uses regression to predict the value of an integer variable $x_{i}$ within the range $[0,1000]$. The maximum possible prediction error for a binary variable, with a range of $[0,1]$, is 1. However, since the range of $x_{i}$ is significantly broader, the maximum prediction error for $x_{i}$ would also be broader. Moreover, additional errors can occur due to rounding that makes the predicted values integers. Therefore, a more accurate ML-based method for integer variables is required. One study (Nair et al. 2020) proposed a method to handle integer variables by representing the values in binary format. To represent $x_{i}$ as a binary sequence, the required sequence length would be $\left\lceil\log_{2}(1000)\right\rceil=10$. Thus, predicting a single decision variable requires multiple dimensions, and the upper/lower bounds of a variable may even be infinite in some cases. This highlights the need for more efficient methods to handle integer decision variables.
To address these limitations, we propose an RL-based method for solving MILP. Figure 1(b) illustrates how the proposed solver derives the optimal solution for an MILP instance. Unlike ML-based primal heuristics which delegate sub-problems to traditional optimizers, the proposed method takes an MILP instance and directly searches for complete feasible solutions. To focus on handling integer variables, we consider a special case of MILP where all variables are integers. Rather than directly predicting the exact values of integer variables, we adopt an indirect approach where the ML model determines whether to increase, decrease, or retain the current value of each decision variable.
@@
To guide these decisions, we design an RL system tailored to MILP, enabling the agent to effectively learn the relationships between decision variables and constraints. The solution search process is divided into two phases: before and after finding the first feasible solution of MILP. We design reward functions for each phase to align with their objectives. During training, the RL agent is updated based on the degree of constraint violations and improvements in the objective value. To capture the relationships among all decision variables effectively, we adopt a Transformer encoder-based GNN as the agent’s architecture. The trained agent processes unseen MILP instances and incrementally improves the solution using the proposed local search strategy. Experimental results demonstrate that the proposed model achieves optimal solutions on a small dataset and finds near-optimal solutions with roughly 1% of the optimal for a larger dataset. Our main contributions are summarized as follows:
@@ 2.3 Graph Neural Networks
(a) Example of a bipartite graph representation for an MILP instance with four variables and three constraints.
@@ 3 Methodology
This section provides a detailed introduction to our RL-MILP solver. Figure 3 demonstrates how the proposed solver incrementally improves feasible solutions. The first step involves transforming the MILP instance in its standard form (Eqs. 1-4) into a bipartite graph. The MILP bipartite graph serves as input to the GNN-based agent. In the second step, the agent trained with an RL algorithm selects actions estimated to yield high rewards. The agent decides for each variable whether to increase, decrease, or retain its value. In the third step, the selected actions are applied to derive a solution for the given instance. The final solution is updated if the new feasible solution $\mathbf{x}$, obtained from the selected actions, is better than the current best feasible solution.
@@ 3.1 Reinforcement Learning for MILP
$\mathcal{R}_{t}\mathrel{+}{=}1,\quad\text{if feasible}\land obj_{t}<obj_{b}$ (8)
@@
where $obj_{r}$ serves as the lower bound that is the objective value of the LP-relaxation of the original MILP problem, and $obj_{b}$ denotes the smallest objective value found so far. The value of $\delta_{b}$, ranging from 0 to 1, indicates how close $obj_{t}$ is to $obj_{b}$ (Eq. 9). Smaller $\delta_{b}$ values indicate greater proximity between the two.
To enhance understanding, we explain the reward function in $phase2$ (Eq. 7) with Figure 5. If a better (i.e., lower) objective value than $obj_{b}$ exists, it must lie between the lower bound $obj_{r}$ and $obj_{b}$. Thus, the agent should explore solutions in this interval to obtain a better $obj_{t}$. Getting closer to $obj_{r}$ improves the $obj$ value, but finding feasible solutions becomes harder in this region because of the tight constraints. Therefore, the agent is encouraged to search for solutions closer to $obj_{b}$, where finding feasible solutions is relatively easier. For this reason, within the interval, higher rewards are given as $obj_{t}$ gets closer to $obj_{b}$ (Case 1 in Eq. 7). Regardless of whether $\textbf{x}_{t}$ is feasible, a reward is given if $obj_{t}$ lies within the interval. In Figure 5, the red and green circles correspond to this first case. We regard the agent’s actions that result in $obj_{t}$ falling outside the interval as incorrect. For infeasible solutions $\textbf{x}_{t}$ outside the interval, a penalty is given proportional to the distance of $obj_{t}$ from $obj_{b}$ (Case 2 in Eq. 7). In Figure 5, the red triangle represents this second case. On the other hand, penalizing feasible solutions outside the interval could undermine the agent’s learning from $phase1$, where the goal is to find feasible solutions. To preserve consistency in encouraging feasible solution discovery across $phase1$ and $phase2$, no penalties are given for feasible solutions outside the interval (Case 3 in Eq. 7). In Figure 5, the green triangle corresponds to this third case. If the agent finds a better feasible solution within the interval, an additional reward of $+1$ is given (Eq. 8), and it corresponds to the green circle in Figure 5. Finally, in both $phase1$ and $phase2$, if the agent does not perform any exploration (i.e., no change (--) for all variables), a penalty of $-10$ is added to $\mathcal{R}_{t}$ to encourage exploration.
@@ 3.2 GNN Architecture of RL Agent
We design the agent’s architecture to solve MILP problems. For the architecture, we adopt a hybrid approach (Rong et al. 2020; Wu et al. 2021) that combines an MPNN-based GNN with a Transformer encoder. Figure 6 illustrates the architecture of the RL agent for solving MILP problems. In our study, the MPNN-based GNN captures local information specific to each decision variable. Meanwhile, the Transformer encoder utilizes the captured local information, along with additional MILP features, as inputs to learn global information across all variables. We encode the following features as inputs to the GNN: the objective value $obj_{t}$, the coefficients of the objective function $\mathbf{c}$, the constraint coefficients $\mathbf{A}$, the right-hand side of the constraints $\mathbf{b}$, the solution values $\mathbf{x}_{t}$, and the bound features $bnd\_lim$ for the variables’ lower and upper bounds.
The MPNN-based GNN aims to capture local information for each decision variable. In an MILP graph, local information can include whether the constraints involving a given decision variable are satisfied. To learn feasibility relationship between variables and constraints, we use $\mathbf{x}_{t}$, $\mathbf{b}$, and $\mathbf{A}$ as the features for variable nodes, constraint nodes, and edges, respectively. Input feature scaling is necessary to ensure stable training. We apply equilibration scaling (Tomlin 1975) to scale $\mathbf{A}$. Equilibration scaling relies on the principle that multiplying each constraint by a non-negative scalar does not change the optimization results of the original MILP instance. We scale $\mathbf{A}$ to the range $[-1,1]$ by dividing each constraint by its own largest absolute coefficient. For $\mathbf{x}_{t}$ and $\mathbf{b}$, whose bounds are difficult to define and thus challenging to scale, we employ a periodic activation-based module, called Periodic Embedding (PE), to embed scalar values into vectors. PE has been shown to improve the performance of DL models in tasks involving numerical feature processing (Gorishniy, Rubachev, and Babenko 2022), such as house price prediction (Pace and Barry 1997) or income prediction (Kohavi et al. 1996). PE is formulated as follows:
@@
Output: Best feasible solution found $\textbf{x}_{b}$
@@ 5 Related Work
A survey paper on ML techniques for solving CO problems (Bengio, Lodi, and Prouvost 2021) categorizes existing studies into three groups. The first group, Learning to Configure Algorithms, leverages ML to optimize the configuration of specific techniques in traditional MILP solvers. For instance, ML can improve solving speed by deciding whether to execute specific operations, such as decomposition (Kruber, Lübbecke, and Parmentier 2017). The second group, Machine Learning Alongside Optimization Algorithms, involves traditional solvers invoking ML repeatedly during the optimization process to support important decisions, such as selecting the next search region (Lodi and Zarpellon 2017). The third group, End-to-End Learning, involves ML directly learning and predicting solutions. This group includes ML-based primal heuristics (Gasse et al. 2019; Nair et al. 2020; Yoon 2022; Han et al. 2023; Cantürk et al. 2024), as well as our method. Existing ML-based primal heuristics output only partial solutions and face challenges in handling non-binary integer variables. Moreover, they rely on supervised learning, which requires high costs for collecting labeled training data. In contrast, our RL-based method does not require labeled data. A previous RL-based study on MILP focuses on heuristics for quickly finding the first feasible solution (Qi, Wang, and Shen 2021) and does not address the search for optimal solutions. To the best of our knowledge, our study is the first to propose an ML-based method that incrementally finds better complete feasible solutions.
@@ 6 Conclusion
In this study, we proposed a novel RL-based MILP solver capable of generating complete feasible solutions, unlike previous approaches that delegate sub-problems to traditional solvers. We designed an RL system that enables the agent to capture the relationships between decision variables and constraints while learning to solve MILP problems. The reward functions are tailored for two distinct phases: before and after finding the first feasible solution. To effectively capture relationships among decision variables, we proposed a Transformer-based GNN architecture for the RL agent. Our experiments demonstrated that the Transformer-based GNN outperformed other variants in terms of RP_Gap. The trained agent achieved optimal solutions for small datasets and near-optimal solutions with roughly 1% of RP_Gap for larger datasets. This study proposes a novel RL-based MILP solver, which establishes a foundation for ML-driven approaches that eliminate reliance on traditional solvers. This work focused on an MILP problem where all variables are integers while future research can extend the proposed method to handle continuous variables as well. Additionally, the local search strategy employed in this study has inherent limitations due to its restricted search region, resulting in challenges to reaching the global optimal solution. To overcome these limitations and avoid local optima, further research on global search strategies for MILP is necessary.

### v2

@@ Abstract
Mixed-Integer Linear Programming (MILP) is an optimization technique widely used in various fields. Existing end-to-end learning methods for MILP generate values for a subset of decision variables and delegate the remaining problem to traditional MILP solvers. However, this approach does not guarantee solution feasibility (i.e., satisfying all constraints) due to inaccurate predictions and primarily focuses on prediction for binary decision variables. When addressing MILP involving non-binary integer variables using machine learning (ML), feasibility issues can become even more pronounced. Since finding an optimal solution requires satisfying all constraints, addressing feasibility is critical. To overcome these limitations, we propose a novel reinforcement learning (RL)-based solver that interacts with MILP to incrementally discover better feasible solutions without relying on traditional solvers. We design reward functions tailored for MILP, which enable the RL agent to learn relationships between decision variables and constraints. Furthermore, we leverage a Transformer encoder-based graph neural network (GNN) to effectively model complex relationships among decision variables. Our experimental results demonstrate that the proposed method can solve MILP problems and find near-optimal solutions without delegating the remainder to traditional solvers. The proposed method provides a meaningful step forward as an initial study in solving MILP problems entirely with ML in an end-to-end manner.
@@ 1 Introduction
The traveling salesman problem (TSP) and the knapsack problem are representative examples of combinatorial optimization (CO) problems that have been extensively studied in operations research and computer science. CO addresses mathematical optimization problems that aim to minimize or maximize the value of a specific objective function. If the objective function and constraints of CO are linear, it is called linear programming (LP). If some decision variables in LP must take integer values, it becomes Mixed-Integer Linear Programming (MILP) (Bengio, Lodi, and Prouvost 2021). MILP is widely used to model various problems (Zhang et al. 2023) such as logistics (Kweon et al. 2024), path planning (Zuo et al. 2020), and energy systems (Ren and Gao 2010).
End-to-end learning refers to the approach in which ML directly learns and predicts solutions for CO problems (Bengio, Lodi, and Prouvost 2021). ML-based primal heuristics are considered as end-to-end learning (Han et al. 2023), which aim to quickly find initial solutions to MILP by reducing the search space (Achterberg, Berthold, and Hendel 2012). Unlike traditional primal heuristics, which rely on expert knowledge and hand-crafted designs (Bengio, Lodi, and Prouvost 2021), the ML-based approach leverages similar patterns shared by MILP instances generated from a specific distribution (Gasse et al. 2022). Figure 1(a) illustrates how ML-based primal heuristics solve MILP. A trained ML model generates a partial solution for a subset of integer variables in a given MILP instance. The partial solution is passed to a traditional optimizer (e.g., Gurobi) that optimizes the remaining sub-problem to derive the final solution.
@@
Figure 1: Solving MILP with end-to-end learning methods.
@@
Existing ML-based primal heuristics have demonstrated their ability to quickly find good solutions by reducing the problem space (Gasse et al. 2019; Nair et al. 2020; Yoon 2022; Han et al. 2023; Cantürk et al. 2024). However, inaccurate ML predictions hinder solution feasibility (i.e., satisfying all constraints), which emphasizes the need to consider feasibility (Han et al. 2023). The infeasibility prevents reaching the optimal solution, posing a significant obstacle to solving MILP. Moreover, most ML-based primal heuristics (Gasse et al. 2019; Yoon 2022; Han et al. 2023; Cantürk et al. 2024) focus solely on predicting values for binary decision variables. However, many real-world problems involve non-binary integer (integer, for short) variables such as logistics (Kweon et al. 2024), maritime transportation (Papageorgiou et al. 2014), and energy systems (Ren and Gao 2010). This highlights the need for techniques capable of effectively handling integer variables.
@@
Infeasibility caused by incorrect ML predictions can be more pronounced for integer variables due to their broader prediction range than binary variables. For instance, imagine a naive ML approach that uses regression to predict the value of an integer variable $x_{i}$ within the range $[0,1000]$. The maximum possible prediction error for a binary variable, with a range of $[0,1]$, is 1. However, since the range of $x_{i}$ is significantly broader, the maximum prediction error for $x_{i}$ would also be broader. Moreover, additional errors can occur due to rounding that makes the predicted values integers. Therefore, a more accurate ML-based method for integer variables is required. Nair et al. have proposed a method to handle integer variables by representing the values in binary format. To represent $x_{i}$ as a binary sequence, the required sequence length would be $\left\lceil\log_{2}(1000)\right\rceil=10$. Thus, predicting a single decision variable requires multiple dimensions, and the upper/lower bounds of a variable may even be infinite in some cases. It calls for more efficient methods to handle integer decision variables.
@@
To address these limitations, we propose an RL-based method for solving MILP. Figure 1(b) illustrates how the proposed solver derives the final solution for an MILP instance. Unlike the existing ML-based methods that focus on predictions for a subset of variables in MILP, the proposed method generates complete feasible solutions. Rather than directly predicting the exact values of integer variables, we adopt an indirect approach where the ML model determines whether to increase, decrease, or retain the current value of each decision variable. To guide these decisions, we design an RL system tailored to MILP, enabling the agent to effectively learn the relationships between decision variables and constraints. The solution search process is divided into two phases: before and after finding the first feasible solution of MILP. We design reward functions for each phase to align with their objectives. During training, the RL agent is updated based on the degree of constraint violations and improvements in the objective value. To capture the relationships among all decision variables effectively, we adopt a Transformer encoder-based GNN as the agent’s architecture. The trained agent processes unseen MILP instances and incrementally improves the solution using the proposed local search strategy. Experimental results demonstrate that the proposed model achieves optimal solutions on a small dataset and finds near-optimal solutions with roughly 1% of the optimal for a larger dataset. Our main contributions are summarized as follows:
@@ 2.3 Graph Neural Networks
(a) Example of a bipartite graph representation for an MILP instance with three variables and two constraints.
@@ 3 Methodology
This section provides a detailed introduction to our RL-MILP solver. Figure 3 demonstrates how the proposed solver incrementally improves feasible solutions. The first step involves transforming the MILP instance in its standard form (Eqs. 1-4) into a bipartite graph. The MILP bipartite graph serves as input to the GNN-based agent. In the second step, the agent trained with an RL algorithm selects actions estimated to yield high rewards. The agent decides for each variable whether to increase, decrease, or retain its value. In the third step, the selected actions are applied to derive a solution for the given instance. The final solution is updated if the new feasible solution $\mathbf{x}$, obtained from the selected actions, is better than the best feasible solution found so far.
@@ 3.1 Reinforcement Learning for MILP
$\mathcal{R}_{t}\mathrel{+}{=}\text{ }C_{\text{reward}},\quad\text{if feasible}\land obj_{t}<obj_{b}$ (8)
@@
where $obj_{r}$ serves as the lower bound that is the objective value of the LP-relaxation of the original MILP problem, and $obj_{b}$ denotes the smallest objective value found so far, $C_{\text{reward}}$ denotes a positive constant which is set to $+1$ in our study. The value of $\delta_{b}$, ranging from 0 to 1, indicates how close $obj_{t}$ is to $obj_{b}$ (Eq. 9). Smaller $\delta_{b}$ values indicate greater proximity between the two.
To enhance understanding, we explain the reward function in $phase2$ (Eq. 7) with Figure 5. If a better (i.e., lower) objective value than $obj_{b}$ exists, it must lie between the lower bound $obj_{r}$ and $obj_{b}$. Thus, the agent should explore solutions in this interval to obtain a better $obj_{t}$. Getting closer to $obj_{r}$ improves the $obj$ value, but finding feasible solutions becomes harder in this region because of the tight constraints. Therefore, the agent is encouraged to search for solutions closer to $obj_{b}$, where finding feasible solutions is relatively easier. For this reason, within the interval, higher rewards are given as $obj_{t}$ gets closer to $obj_{b}$ (Case 1 in Eq. 7). Regardless of whether $\textbf{x}_{t}$ is feasible, a reward is given if $obj_{t}$ lies within the interval. In Figure 5, the red and green circles correspond to this first case. We regard the agent’s actions that result in $obj_{t}$ falling outside the interval as incorrect. For infeasible solutions $\textbf{x}_{t}$ outside the interval, a penalty is given proportional to the distance of $obj_{t}$ from $obj_{b}$ (Case 2 in Eq. 7). In Figure 5, the red triangle represents this second case. On the other hand, penalizing feasible solutions outside the interval could undermine the agent’s learning from $phase1$, where the goal is to find feasible solutions. To preserve consistency in encouraging feasible solution discovery across $phase1$ and $phase2$, no penalties are given for feasible solutions outside the interval (Case 3 in Eq. 7). In Figure 5, the green triangle corresponds to this third case. If the agent finds a better feasible solution within the interval, an additional reward of $C_{\text{reward}}$ is given (Eq. 8), and it corresponds to the green circle in Figure 5. Finally, in both $phase1$ and $phase2$, if the agent does not perform any exploration (i.e., no change (--) for all variables), a penalty of $-10$ is added to $\mathcal{R}_{t}$ to encourage exploration.
@@ 3.2 GNN Architecture of RL Agent
For the agent’s architecture, we adopt a hybrid approach that combines an MPNN-based GNN with a Transformer encoder (Rong et al. 2020; Wu et al. 2021). Figure 6 illustrates the architecture of the RL agent for solving MILP problems. In our study, the MPNN-based GNN captures local information specific to each decision variable. Meanwhile, the Transformer encoder utilizes the captured local information, along with additional MILP features, as inputs to learn global information across all variables. We encode the following features as inputs to the GNN: the objective value $obj_{t}$, the coefficients of the objective function $\mathbf{c}$, the constraint coefficients $\mathbf{A}$, the right-hand side of the constraints $\mathbf{b}$, the solution values $\mathbf{x}_{t}$, and the bound features $bnd\_lim$ for the variables’ lower and upper bounds.
The MPNN-based GNN aims to capture local information for each decision variable. In an MILP graph, local information can include whether the constraints involving a given decision variable are satisfied. To learn feasibility relationship between variables and constraints, we use $\mathbf{x}_{t}$, $\mathbf{b}$, and $\mathbf{A}$ as the features for variable nodes, constraint nodes, and edges, respectively. Input feature scaling is necessary to ensure stable training. We apply equilibration scaling (Tomlin 1975) to scale $\mathbf{A}$. Equilibration scaling relies on the principle that multiplying each constraint by a non-negative scalar does not change the optimization results of the original MILP instance. We scale $\mathbf{A}$ to the range $[-1,1]$ by dividing each constraint by its own largest absolute coefficient. For $\mathbf{x}_{t}$ and $\mathbf{b}$, whose bounds are difficult to define and thus challenging to scale, we employ a periodic activation-based module, called Periodic Embedding (PE), to embed scalar values into vectors. PE has been shown to improve the performance of DL models in tasks involving numerical feature processing (Gorishniy, Rubachev, and Babenko 2022), such as house price prediction (Pace and Barry 1997) and income prediction (Kohavi et al. 1996). PE is formulated as follows:
@@
Output: Best feasible solution $\textbf{x}_{b}$
@@ 5 Related Work
Bengio, Lodi, and Prouvost have categorized ML techniques for solving CO problems into three groups. The first group, Learning to Configure Algorithms, leverages ML to optimize the configuration of specific techniques in traditional MILP solvers. For instance, ML can improve solving speed by deciding whether to execute specific operations, such as decomposition (Kruber, Lübbecke, and Parmentier 2017). The second group, Machine Learning Alongside Optimization Algorithms, involves traditional solvers invoking ML repeatedly during the optimization process to support important decisions, such as selecting the next search region (Lodi and Zarpellon 2017). The third group, End-to-End Learning, involves ML directly learning and predicting solutions. This group includes ML-based primal heuristics (Gasse et al. 2019; Nair et al. 2020; Yoon 2022; Han et al. 2023; Cantürk et al. 2024), as well as our method. Existing ML-based primal heuristics output only partial solutions and face challenges in handling non-binary integer variables. Moreover, they rely on supervised learning, which requires high costs for collecting labeled training data. In contrast, our RL-based method does not require labeled data. A previous RL-based study on MILP focuses on heuristics for quickly finding the first feasible solution (Qi, Wang, and Shen 2021) and does not address the search for optimal solutions. To the best of our knowledge, our study is the first to propose an ML-based method that incrementally finds better complete feasible solutions.
@@ 6 Conclusion
In this study, we proposed a novel RL-based MILP solver capable of generating complete feasible solutions, unlike previous approaches that delegate sub-problems to traditional solvers. We designed an RL system that enables the agent to capture the relationships between decision variables and constraints while learning to solve MILP problems. The reward functions are tailored for two distinct phases: before and after finding the first feasible solution. To effectively capture relationships among decision variables, we proposed a Transformer-based GNN architecture for the RL agent. Our experiments demonstrated that the Transformer-based GNN outperformed other variants in terms of RP_Gap. The trained agent achieved optimal solutions for small datasets and near-optimal solutions with roughly 1% of RP_Gap for larger datasets. This study proposes a novel RL-based MILP solver, which establishes a foundation for ML-driven approaches that eliminate reliance on traditional solvers. This work focused on an MILP problem where all variables are integers while future research can extend the proposed method to handle continuous variables as well. In addition, the local search strategy faces challenges in reaching the global optimal solution due to its restricted search region. To overcome this limitation, further research on search strategies for MILP is necessary. Building on our study as a cornerstone, future work could explore ML approaches that quickly search for high-quality solutions for large-scale problems, where traditional solvers may struggle due to their exponential time complexity.